# Debiased machine learning for combining probability and non-probability survey data

### v1

@@ 10 Discussion
We have described a debiased machine learning approach for estimating the mean of $Y$ from a nonprobability sample when auxiliary information on $X$ are available from a probability sample. This method yields an asymptotically normally distributed estimator with a standard error that can estimated as easily as can the standard error of a Horwitz-Thompson estimator. To prove this, we have relied on cross-fitting and have (unless cluster sampling probabilities are all equal) discarded some information when estimating the nuisance function $\hat{\pi}^{B}$. For the closely related problem of estimating $E(Y)$ for an infinite population using iid data, an alternative to cross-fitting is to rely on the Donsker condition [20]. This limits the choice of data-adaptive estimators of $P(R=1\mid X)$ and $E(Y\mid X)$ to those that satisfy the Donsker condition. Nevertheless, it would be interesting to establish whether the Donsker condition would suffice in the setting we study in this article.
@@
Some researchers have previously used DR estimators with data-adaptive estimation of nuisance functions, e.g. [3] and [11]. However, they did not establish the validity of this approach. Indeed [11] comment (page 276) that “there are hardly any results on the theoretical properties of variance estimators when more complex machine learning models and techniques are used.” Moreover, these researchers used data-adaptive estimators that do not necessarily satisfy the Donsker condition and without using cross-fitting, which is potentially problematic even in the simpler situation of estimating $E(Y)$ for an infinite population using iid data.
@@
To establish asymptotic properties of our estimators, we have assumed that the finite population is generated from a particular superpopulation model, so that it grows in a particular way. It would be interesting to investigate whether asymptotic properties can be established when the finite population is allowed to grow in another way. We also used subsampling of an ‘active subset’ of clusters when calculating $\hat{\pi}^{B}$. This should have no effect on asymptotic efficiency, and in our simulation study we observed no loss of efficiency from doing it. However, if loss of efficiency in finite samples is a concern, it could be mitigated by using repeated cross-fitting [8]. As with other debiased machine learning estimators using cross-fitting, repeated cross-fitting would also have the advantage of reducing the finite-sample Monte Carlo variation induced by the randomness of the process of partitioning units (in our case, clusters) among folds, although at the cost of increasing computation time.
@@
The approach we have presented is for probability samples that do not use stratified sampling. It is straightforward in principle to generalise it to deal with stratified sampling. Now the assignment of clusters to the $K$ folds should be done within each stratum. This should be fairly unproblematic when the probability sample uses simple random sampling of clusters without replacement within each stratum or when the number of clusters sampled from each stratum is large. When, however, the cluster sampling probabilities within strata vary and the number of sampled clusters within strata is small, there may be two problems. First, the subsampling of active subsets of clusters could lead to substantial loss of information for estimating $\hat{\pi}^{B}$ in this situation. Second, even with the subsampling of an active subset, the asymptotic property that $\hat{\pi}^{B}_{k}$ is uninformative about $R^{A}$ values in fold $k$ may be a poor approximation in this situation. The scenario of sampling few clusters per stratum is a problem worthy of further research.
@@
Finally, several researchers have considered the problem of efficiently combining an IPW or DR estimator of the mean of $Y$ with a Horwitz-Thompson or Hajek estimator of this same mean based on Sample A alone. Gao and Yang [18] consider the DR estimator $\hat{\theta}_{1}$ and use parametric nuisance models. Seaman et al. (2025) discuss this approach and extend it to $\hat{\theta}_{{\rm CLW}2}$ and IPW estimators, again using parametric nuisance models. Rueda et al. (2023)[13] used an IPW estimator with random forest or XGboost to estimate the nuisance function. Since, when Condition C1 is satisfied, the debiased machine learning estimators of $\bar{Y}$ that we have presented here have the same asymptotic distribution as the corresponding DR estimators of $\bar{Y}$ that use two correctly specified parametric nuisance models, the combining methods discussed by Seaman et al. (2025) are equally appropriate.
@@ Acknowledgements
A1 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for simple random sampling of clusters without replacement
@@
Now consider the second variance in expression (A4). As we argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, for any fixed function $\hat{m}_{k}$,
@@
is the variance of the Horwitz-Thompson estimator of the population mean of $\hat{m}_{k}(X)-m_{0}(X)$ in individuals in $\mathcal{S}_{k}$ given $\mathcal{F}_{J}$ when the clusters in $\mathcal{S}_{k}$ are selected by simple random sampling without replacement and individuals within clusters are sampled according to the original second-stage sampling procedure. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020 [4]).
@@
As argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, $E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=\pi^{A}_{i}$. Hence, expression (A7) reduces to
@@
A2 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for random sampling of clusters without replacement with unequal probabilities
@@
Consider the empirical process term for fold $k$. Equations (A2) and (A3) still apply when clusters are sampled without replacement with unequal probabilities. Consider the variance in expression (A3).
@@
By a very similar argument to that used in Section 4, we see that, because of the way that fold $S_{k}$ has been chosen, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}$ corresponds to stratified simple random sampling of $M^{(l)}_{k}$ clusters without replacement from the stratum of $J^{(l)}_{k}$ clusters with the same value of $\pi^{C(l)}$ followed by the original second-stage sampling of individuals within clusters. The probability of sampling individual $i$ is the original probability $\pi^{A}_{i}$ multiplied by $M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})$, because the fraction of clusters sampled from this stratum is $M^{(l)}_{k}/J^{(l)}_{k}$, rather than $\pi^{C(l)}$. Hence, for any fixed function $\hat{m}_{k}$, expression (A12) is the variance of the Horwitz-Thompson estimator of the population mean in the $k$th fold of $\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\{M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$ when clusters are sampled in this stratified way. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020 [4]). Consequently, if $\hat{m}_{k}\xrightarrow{p}m_{0}$ as $M\rightarrow\infty$, then this variance is $o_{p}(M^{-1})$. Hence, the conditional expectation given $\mathcal{F}_{J},S_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ of this variance is also $o_{p}(M^{-1})$. Therefore, the first of the two terms in expression (A10) is $o_{p}(M^{-1})$, as required.
@@
Equation (A7) still applies when clusters are sampled without replacement with unequal probabilities. As before, because of the way that the folds have been chosen, and using equation (A14), we have
@@
The final term in expression (A21) is $O_{p}(M^{-1/2})$ for fixed $\hat{m}_{k}$. If $\hat{m}_{k}\xrightarrow{P}m_{0}$, then this term becomes $o_{p}(M^{-1/2})$, and so can be ignored. The rest of expression (A21) is the same as expression (A8). Hence, we are back to the situation we have where sampling of clusters is by simple random sampling without replacement. Thus, under the conditions given there, the remainder term is $o_{p}(M^{-1/2})$, as required.

### v2

@@ 10 Discussion
We have described a debiased machine learning approach for estimating the mean of $Y$ from a nonprobability sample when auxiliary information on $X$ are available from a probability sample. This method yields an asymptotically normally distributed estimator with a standard error that can estimated as easily as can the standard error of a Horwitz-Thompson estimator. To prove this, we have relied on cross-fitting. For the closely related problem of estimating $E(Y)$ for an infinite population using iid data, an alternative to cross-fitting is to rely on the Donsker condition[21]. This limits the choice of data-adaptive estimators of $P(R=1\mid X)$ and $E(Y\mid X)$ to those that satisfy the Donsker condition. Nevertheless, it would be interesting to establish whether the Donsker condition would suffice in the setting we study in this article.
@@
Some researchers have previously used DR estimators with data-adaptive estimation of nuisance functions for combining probability and non-probability survey data, e.g. [3] and [12]. However, they did not establish the validity of this approach. Indeed [12] comment (page 276) that “there are hardly any results on the theoretical properties of variance estimators when more complex machine learning models and techniques are used.” Moreover, these researchers used data-adaptive estimators that do not necessarily satisfy the Donsker condition and without using cross-fitting, which is potentially problematic even in the simpler situation of estimating $E(Y)$ for an infinite population using iid data.
@@
In general, we have used subsampling of an ‘active subset’ of clusters when calculating $\hat{\pi}^{B}$. This should have no effect on asymptotic efficiency, and in our simulation study we observed no loss of efficiency from doing it. However, if loss of efficiency in finite samples is a concern, it could be mitigated by using repeated cross-fitting[8]. As with other debiased machine learning estimators using cross-fitting, repeated cross-fitting would also have the advantage of reducing the finite-sample Monte Carlo variation induced by the randomness of the process of partitioning units (in our case, clusters) among folds, although at the cost of increasing computation time.
As discussed in Section 4, if SRSWOR is used to select the $M$ clusters for Sample A and $M$ is an integer multiple of the number of folds $K$, there is no need to subsample an active subset. Appendix A11 shows this is true also when the clusters are sampled using conditional Poisson sampling.
@@
The approach we have presented is for probability samples that do not use stratified sampling. It is straightforward in principle to generalise it to deal with stratified sampling. Now the assignment of clusters to the $K$ folds would be done within each stratum. This should be fairly unproblematic when the probability sample uses SRSWOR of clusters within each stratum or when the number of clusters sampled from each stratum is large. When, however, the cluster sampling probabilities within strata vary and the number of sampled clusters within strata is small, there may be two problems. First, the subsampling of active subsets of clusters could lead to substantial loss of information for estimating $\hat{\pi}^{B}$ in this situation. Second, even with the subsampling of an active subset, the asymptotic property that $\hat{\pi}^{B}_{k}$ is uninformative about $R^{A}$ values in fold $k$ may be a poor approximation in this situation. The scenario of sampling few clusters per stratum is a problem worthy of further research.
To establish asymptotic properties of our estimators, we have assumed that the finite population is generated from a particular superpopulation model, so that it grows in a particular way. It would be interesting to investigate whether asymptotic properties can be established when the finite population is allowed to grow in another way.
@@
Finally, several researchers have considered the problem of efficiently combining an IPW or DR estimator of the mean of $Y$ with a Horwitz-Thompson or Hajek estimator of this same mean based on Sample A alone when $Y$ is observed in both samples. Gao and Yang (2023)[19] consider the DR estimator $\hat{\theta}_{1}$ and use parametric nuisance models. Seaman et al. (2025)[26] discuss this approach and extend it to $\hat{\theta}_{{\rm CLW}2}$ and IPW estimators, again using parametric nuisance models. Rueda et al. (2023)[14] used an IPW estimator with random forest or XGBoost to estimate the nuisance function. Since, when Condition C1 is satisfied, the debiased machine learning estimators of $\bar{Y}$ that we have presented here have the same asymptotic distribution as the corresponding DR estimators of $\bar{Y}$ that use two correctly specified parametric nuisance models, the combining methods discussed by Seaman et al. (2025) should apply.
@@ Acknowledgements
A1 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for SRSWOR of clusters
@@
Now consider the second variance in expression (A4). As we argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, for any fixed function $\hat{m}_{k}$,
@@
is the variance of the Horwitz-Thompson estimator of the population mean of $\hat{m}_{k}(X)-m_{0}(X)$ in individuals in $\mathcal{S}_{k}$ given $\mathcal{F}_{J}$ when the clusters in $\mathcal{S}_{k}$ are selected by SRSWOR and individuals within clusters are sampled according to the original second-stage sampling procedure. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020[4]).
@@
As argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, $E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=\pi^{A}_{i}$. Hence, expression (A7) reduces to
@@
A2 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for SWOR of clusters with unequal probabilities
@@
Consider the empirical process term for fold $k$. Equations (A2) and (A3) still apply when clusters are selected by SWOR with unequal probabilities. Consider the variance in expression (A3).
@@
By a very similar argument to that used in Section 4, we see that, because of the way that fold $S_{k}$ has been chosen, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}$ corresponds to stratified SRSWOR of $M^{(l)}_{k}$ clusters from the stratum of $J^{(l)}_{k}$ clusters with the same value of $\pi^{C(l)}$ followed by the original second-stage sampling of individuals within clusters. The probability of sampling individual $i$ is the original probability $\pi^{A}_{i}$ multiplied by $M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})$, because the fraction of clusters sampled from this stratum is $M^{(l)}_{k}/J^{(l)}_{k}$, rather than $\pi^{C(l)}$. Hence, for any fixed function $\hat{m}_{k}$, expression (A12) is the variance of the Horwitz-Thompson estimator of the population mean in the $k$th fold of $\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\{M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$ when clusters are sampled in this stratified way. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020[4]). Consequently, if $\hat{m}_{k}\xrightarrow{p}m_{0}$ as $M\rightarrow\infty$, then this variance is $o_{p}(M^{-1})$. Hence, the conditional expectation given $\mathcal{F}_{J},S_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ of this variance is also $o_{p}(M^{-1})$. Therefore, the first of the two terms in expression (A10) is $o_{p}(M^{-1})$, as required.
@@
Equation (A7) still applies when clusters are selected using SWOR with unequal probabilities. As before, because of the way that the folds have been chosen, and using equation (A14), we have
@@
The final term in expression (A21) is $O_{p}(M^{-1/2})$ for fixed $\hat{m}_{k}$. If $\hat{m}_{k}\xrightarrow{P}m_{0}$, then this term becomes $o_{p}(M^{-1/2})$, and so can be ignored. The rest of expression (A21) is the same as expression (A8). Hence, we are back to the situation we have where clusters are selected by SRSWOR. Thus, under the conditions given there, the remainder term is $o_{p}(M^{-1/2})$, as required.